OpenStudy (breanna1999):

The function f(x) = 4(2)x represents the growth of a butterfly population every year in a remote swamp. Jan wants to manipulate the formula to an equivalent form that calculates five times a year, not just once a year. Which function is correct for Jan's purpose, and what is the new growth rate?

OpenStudy (therawbugeyes):

ok this i can help u on

OpenStudy (breanna1999):

would it be f(x)= 4(2^1/5)^5?

OpenStudy (therawbugeyes):

i do need you answer choices tho

OpenStudy (breanna1999):

f(x) = 4(1.15)x; growth rate is 5% f(x) = 4(1.15)5x; growth rate is 15% f(x) = 4(2)x; growth rate is 200% f(x) = 4(2)x, growth rate is 5%

OpenStudy (therawbugeyes):

is this algebra from flvs

OpenStudy (breanna1999):

yes. Algebra 2 segment review

OpenStudy (therawbugeyes):

ok yes it would be what you picked i need some help also can you see if you can help

OpenStudy (breanna1999):

With what??

OpenStudy (therawbugeyes):

Question 12 (Multiple Choice Worth 1 points) (03.03 MC) The equation below represents Function A and the graph represents Function B: Function A f(x) = 6x – 1 Function B graph of line going through ordered pairs 1, 4 and negative 1, negative 2 and negative 2, negative 5 Which equation best compares the slopes of the two functions? Slope of Function B = 2 x Slope of Function A. Slope of Function A = Slope of Function B Slope of Function A = 2 x Slope of Function B

OpenStudy (therawbugeyes):

this is also algerbra from flvs

OpenStudy (breanna1999):

Id the graph line going up or down?

OpenStudy (breanna1999):

oh its the last one on the list slope a= 2x slope b

OpenStudy (therawbugeyes):

theres the graph

OpenStudy (therawbugeyes):

so it is D

OpenStudy (breanna1999):

Slope of function a= 2x slope of function b is your answer

OpenStudy (therawbugeyes):

ok i have one more please i will medal you if you can help

OpenStudy (breanna1999):

okay go for it

OpenStudy (therawbugeyes):

thxs

OpenStudy (therawbugeyes):

The graph below plots a function f(x): graph of line segment going through ordered pairs 0,150 and 3, 0 If x represents time, the average rate of change of the function f(x) in the first three seconds is ___. Answer for Blank 1:

OpenStudy (therawbugeyes):

and theres the graph

OpenStudy (breanna1999):

answer choices?

OpenStudy (breanna1999):

i got -50 as the average rate of change

OpenStudy (therawbugeyes):

there are none i have to type the answer thats also what i got

OpenStudy (therawbugeyes):

do you have anymore you need help with

OpenStudy (breanna1999):

Not at the moment. Thank you :)

OpenStudy (therawbugeyes):

ok well im staying here if you need help lleave this opean ok

OpenStudy (breanna1999):

okays

OpenStudy (therawbugeyes):

hey so i have another and this one is a bit tricky i hate some of the flvs test cuz they play on words:(

OpenStudy (breanna1999):

i can try and help

OpenStudy (therawbugeyes):

ok awsome thank you so much how do i fan you btw

OpenStudy (therawbugeyes):

Choose the equation of the horizontal line that passes through the point (−5, 1). y = −5 x = 1 y = 1 x = −5

OpenStudy (breanna1999):

y= horizontal line so y=1

OpenStudy (therawbugeyes):

ok thank you so much

OpenStudy (therawbugeyes):

do you need anyhelp

OpenStudy (therawbugeyes):

do you need anyhelp

OpenStudy (breanna1999):

What is the sum of the arithmetic sequence 3, 9, 15..., if there are 26 terms? 1,452 1,728 2,028 2,268

OpenStudy (therawbugeyes):

ok 1 sec

OpenStudy (therawbugeyes):

the answer should be A

OpenStudy (therawbugeyes):

ok

OpenStudy (therawbugeyes):

Anymore?

OpenStudy (breanna1999):

not at the moment

OpenStudy (breanna1999):

What is the solution of log3x − 12 729 = 6? x = −5 x = −3 x = 3 x = 5

OpenStudy (therawbugeyes):

i dont even know that one sorry

OpenStudy (breanna1999):

it was 5

OpenStudy (therawbugeyes):

ohh sorry

OpenStudy (therawbugeyes):

If f(x) = 2x2 − 10, find f(5). 90 10 7.5 40 @breanna1999

OpenStudy (breanna1999):

10? do you have a picture of it